OpenStudy (anonymous):

the square root of "3x+7=x-1 step by step please

OpenStudy (anonymous):

\[\sqrt{3x+7}=x-1\]

OpenStudy (anonymous):

yay your back lol

OpenStudy (anonymous):

Okay, same as last time but actually a bit easier. Square both sides as the first step.

OpenStudy (anonymous):

yes done that. my answers are incorrect though x=3,-2 dont work.

OpenStudy (anonymous):

Okay, so after you square both sides you got \(3x+7=(x-1)^2\), right? Then expand the right side to \(3x+7=x^2-2x+1\)? With me so far?

OpenStudy (anonymous):

yes got that.

OpenStudy (anonymous):

So \(0=x^2-5x-6\), which is \(0=(x-6)(x+1)\)

OpenStudy (anonymous):

oh lol i did 0=(x-3)(x+2)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

Yeah, that would have been \(0=x^2-x-6\). When in doubt, multiply it back out.

OpenStudy (anonymous):

lol i always make small mistake that i cant figure out why the problem turns out wrong last problem i just added the 1 instead of subtracting it and couldnt figure that out until you told me lol thanks so much for your help!! im going to become your fan =p

OpenStudy (anonymous):

Haha any time